Beginning accounts receivable 164,000
Beginning allowance for doubtful accounts 5,400
Sales 2,000,000
Collections on account 1,720,000
Accounts written off 11,800
Collections of accounts previously written off 1,100
Current balance of bad debt expense 0
Uncollectible accounts as a percentage of receivables 4.00%

Using the balance sheet approach, what is the ending balance of bad debt expense?
Answer #1

Ending balance of account receivable = 164000+2000000-1720000-11800 = 432200

Ending balance of bad debt expense = 432200*4% = 17288